Paul's Peanuts has total fixed costs of $23,996. If the​ company's contribution margin is 20​%, the income tax rate is 35​% and the selling price of a box of Peanuts is $20​, how many boxes of Peanuts would the company need to sell to produce a net income of $13,975​?

A.10,750
B.1,680
C.11,374
D. 2,844
Answer #1

Contribution margin per unit = $20*20% = $4

Let no. of units sold be A

[(Contribution margin per unit x no. of units sold) - Fixed cost] x (1-0.35) = Net income

[(4* A) - 23996]*(1-0.35)] = 13975

No. of units sold = 13975/4 = 11374

Option C. is correct answer.